# Spin Photonic Topological Insulator Leaky Wave Antenna By Armchair Arrangement

### v4

@@
Spin Photonic Topological Insulator Leaky Wave Antenna By Armchair Arrangement
††journal: opticajournal††articletype: Research ArticleA new X-band leaky wave antenna has been developed using spin photonic topological insulators. This antenna features a hexagonal unit cell arranged in an armchair configuration. This arrangement provides advantages over the zigzag configuration, particularly by offering a wider leaky region and a more suitable pattern. To design the structure, a parametric study on the cell dimensions has been conducted, and another study has designed the transition region to couple the topological structure with the classical line. The proposed antenna has a low profile and illuminates two sides of the structure simultaneously. Additionally, it offers a 53-degree scanning range and a bandwidth of 2.7 GHz, making it a groundbreaking improvement over other leaky wave antennas that utilize photonic topological insulators. The proposed antenna is an unbalanced Composite Right/Left-Handed (CRLH) leaky wave antenna capable of radiating in both backward and forward directions. Notably, as it transitions through the broadside within its scanning range, there is no significant drop in performance, even in the presence of an open stop band. To the best of our knowledge, this characteristic is unique among unbalanced CRLH leaky wave antennas.
1 Introduction
@@
For guiding applications, the modes of the structure should be situated in the slow wave region. Conversely, for radiating applications, the propagating modes need to be positioned in the fast wave region[10]. The realized leaky wave antennas (LWAs) using metasurfaces in 1D[11, 12, 13] and 2D[14, 15, 16] dimensions have demonstrated versatility in achieving low-profile multi-beam configurations[17], arbitrary shapes[18], and wide scanning ranges[19]. A new application of metasurfaces enables propagation along a line between two impedance surfaces that support TE/TM polarization, which is referred to as line waves[20, 21, 22, 23, 24]. These structures offer an ultra-wide operating bandwidth; however, they are prone to scattering when encountering sharp turns or structural defects[25].
Photonic topological insulators (PTIs) are a recent development designed for use in guiding-wave and antenna applications[26]. They are robust against scattering caused by manufacturing defects and sharp turns[27]. One type of PTI is known as spin PTIs, which preserve time-reversal symmetry and are realized using spin-degenerate metamaterials[28]. Spin PTI metasurfaces are introduced using a complementary hexagonal unit cell[29], as illustrated in Fig. 1. This unit cell exhibits fourfold degeneracy at the symmetry points of K/K’ in its irreducible Brillouin zone (IRBZ). By combining the complementary parts, strong effective magnetoelectric coupling results in a nontrivial band gap at the K/K’ points. The topological structure is formed by aligning the unit cell with its flipped version, creating edge modes at the interface. The unidirectional propagation of line waves at the interface of two structural components is investigated in zig-zag and armchair configurations.
Using photonic topological insulators (PTIs) for leaky wave antenna applications is not limited to spin PTIs[30, 31]. However, such structures require YIG photonic crystals and a static magnetic field bias, which limits their practicality. In contrast, there is no need for such requirements in spin PTI metasurfaces.
A leaky wave antenna based on spin PTI and utilizing a zigzag arrangement has been investigated in [32]. However, this approach presents several significant challenges, including a narrow bandwidth, limited backward radiation scanning, coupling between radiating elements, and an inadequate radiation pattern (see Fig.1).
@@
This paper proposes that adopting an armchair arrangement, instead of a zigzag design, effectively addresses the identified issues and facilitates backward-to-forward beam scanning. A crucial point is that broadside radiation in unbalanced CRLH leaky wave antennas presents significant challenges, often leading to a sudden drop in gain due to an open stopband. However, in the proposed antenna with the armchair arrangement, the impact of this phenomenon is significantly reduced. Simulations were conducted using Ansys HFSS and were subsequently repeated in CST Studio to validate the results.
2 Motivation to use armchair arrangement
The concept of utilizing the armchair arrangement as a leaky wave antenna derives directly from the trapezoidal serpent line microstrip used in antenna applications[34]. As illustrated in Fig.2, there is a clear analogy between these two structures. Initially, to compare the zigzag and armchair configurations, the dimensions were selected to match those in [32]. Specifically, a periodic length of 7 mm and a border width of 0.25 mm were used. The substrate employed is Rogers/Duroid 5880, which has a thickness of 0.127 mm, a permittivity of 2.2, and a tangent loss factor of 0.0009. It is important to note that these dimensions are not final and may vary depending on the specific application. The parametric study on period length and border width identifies the dimensions necessary for designing an X-band leaky wave.
@@ 2 Motivation to use armchair arrangement
2.1 Parametric Study
The objective was to design an X-band leaky wave antenna using a spin topological insulator featuring a hexagonal lattice arranged in an armchair configuration. Therefore, it is essential to analyze the cell parameter to determine the appropriate cell periodicity and border width. The Fig.4 illustrates how the limits of the bandgap vary with changes in the period length and border width. Specifically, Fig. 4 demonstrates that both the maximum and minimum bandgap will decrease as the period length increases. Furthermore, the decline rate of the upper limit (the minimum of the second degenerate band) is greater than that of the lower limit (the maximum of the first degenerate band). On the other hand, Fig.4 shows that the increasing of the border width will lead to a rise in the bandgap width, with a more significant rate of increase observed for shorter period lengths.
@@ 2.1 Parametric Study
2.2 Matching ASL with Topological structure
@@
The field profile of the ASL and the identified cut planes are presented in Fig. 7, where the red and blue areas demonstrate the highest and the lowest amplitude of electric field,respectively. The direction of the triangles shows the direction of tangential electric field to the cut plane. All of these profiles exhibit confinement of the electric field in the interface of two side of structure but some of them have more similarity to the electric field of ASL cross section (see 7). Among these field profiles, those displayed in Figs. 7, 7 show lesser similarity to the ASL profile in Fig. 7 than the others, Because those cut lines are not perpendicular to the propagation direction.
@@
3 Antenna application
The main advantage of using the armchair arrangement in antenna applications is that it effectively places most of the topological bandgap within the fast wave region. According to the findings in Appendix A, this design enables the proposed structure to scan through the broadside, a capability that conventional unbalanced CRLH leaky wave antennas do not offer.
@@ 3 Antenna application
The variation of maximum realized gain with increasing frequency, as analyzed using Ansys HFSS and CST Studio, is illustrated in Fig. 14. The results are consistent and provide further validation of the simulations. The slight decline observed in the frequency range of 9.5 to 10 GHz is attributed to the inherent bandgap of the armchair arrangement. In contrast, to the best of our knowledge, other unbalanced CRLH leaky wave antennas do not exhibit this characteristic [37, 2, 38, 39]. Another noteworthy feature is the nearly constant realized gain across most of the operational frequency range, which typically ranges from 4 to 6 dB.
4 conclusion
A leaky wave antenna is proposed using spin photonic topological insulators (PTIs). It features a hexagonal complementary unit cell arranged in an armchair configuration. The primary advantage of using the armchair arrangement in leaky wave antenna applications is that most of topological bandgap is positioned within the fast wave region. In contrast, for the zigzag arrangement, most of the topological bandgap falls within the slow wave region. The design of the proposed antenna allows for a scanning range of 53 degrees, from 126 to 73 degrees, within an approximate bandwidth of 2.7 GHz. Importantly, the antenna sustains its performance at broadside angles without experiencing a significant decrease, even in the presence of an open stopband.
In Table 1, a comparison of the proposed leaky wave antennas utilizing topological insulators is presented. The table indicates that the antenna designed with an armchair arrangement offers the widest bandwidth and the highest scanning range. The HPMW of the proposed antenna is a weaker point compared to other references, particularly [33] and [30]. This limitation can be improved by increasing the antenna’s length. However, none of the references listed in the table are capable of radiating in a broadside direction, except for [30], which has significant drawbacks. These include only having a single beam, a narrow bandwidth, a bulky design, and the need for a static magnetic field bias. In contrast, the proposed antenna not only radiates broadside but also offers two beams, a wide bandwidth, a low-profile design, and does not require any additional facilities.
@@
Appendix : The Zig-Zag and Armchair Arrangements Integration
@@
The realized gain pattern at 15.8 GHz, where radiation occurs in a broadside direction, is shown in Fig. 17. A comparison between the patterns of the zigzag arrangement (Fig. 1) and the armchair arrangement reveals that the armchair structure is more suitable.
@@
At 15.8 GHz, the realized gain in the open stopband is less than -3 dB compared to the values observed at the ends of the scanning range. To the best of our knowledge, this unusual phenomenon does not occur in other unbalanced CRLH leaky wave antennas. The 36-degree scan at both the top and bottom of the antenna surpasses the findings reported in [33], matching those in [32], even as it crosses the broadside.
Funding This research received no specific grant from any funding agency in the public, commercial, or not-for-profit sectors.
Acknowledgment We would like to thank Prof. A. Zeidaabadi Nezhad form Isfahan University of Technology and Dr. R.J.B. Davis from University of California, San Diego for the fruitful and illuminating discussions.
Disclosures The authors declare no conflicts of interest.
Data Availability The data that support the findings of this study are available from the corresponding author upon reasonable request.

### v5

# Unbalanced CRLH Leaky-wave Antenna With Broadside Radiation Based On Spin Photonic Topological Insulator Featured Hexagonal Configuration In Armchair Arrangement

@@
Abstract
A new X-band leaky-wave antenna has been developed using spin photonic topological insulators. This antenna features a hexagonal unit cell arranged in an armchair configuration. This arrangement provides advantages over the zigzag configuration, particularly by offering a wider operational region and a more suitable pattern. To design the structure, a parametric study on the cell dimensions has been conducted, and another study has designed the transition region to couple the topological structure with the classical line. The proposed antenna has a low profile and illuminates two sides of the structure simultaneously. Additionally, it offers a 53-degree scanning range and a bandwidth of 2.7 GHz, making it a groundbreaking improvement over other leaky-wave antennas that utilize photonic topological insulators. The proposed antenna is an unbalanced CRLH leaky-wave antenna capable of radiating in both backward and forward directions. Notably, as it transitions through the broadside within its scanning range, there is no significant drop in performance, even in the presence of an open stop band. To the best of our knowledge, this characteristic is unique among unbalanced CRLH leaky-wave antennas.
Index Terms:
Leaky-wave Antenna, CRLH , Topological Metasurface, Armchair Configuration=-21pt
I Introduction
@@
For guiding applications, the modes of the structure should be situated in the slow wave region. Conversely, for radiating applications, the propagating modes need to be positioned in the fast wave region[10]. The realized leaky-wave antennas (LWAs) using metasurfaces in 1D[11, 12, 13] and 2D[14, 15, 16] dimensions have demonstrated versatility in achieving low-profile multi-beam configurations[17], arbitrary shapes[18], and wide scanning ranges[19]. A new application of metasurfaces enables propagation along a line between two impedance surfaces that support TE/TM polarization, which is referred to as line waves[20, 21, 22, 23, 24]. These structures offer an ultra-wide operating bandwidth; however, they are prone to scattering when encountering sharp turns or structural defects[25].
Photonic topological insulators (PTIs) are a recent development designed for use in guiding-wave and antenna applications[26]. They are robust against scattering caused by manufacturing defects and sharp turns[27]. One type of PTI is known as spin PTIs, which preserve time-reversal symmetry and are realized using spin-degenerate metamaterials[28]. Spin PTI metasurfaces are introduced using a complementary hexagonal unit cell[29], as illustrated in Fig. LABEL:HexCell. This unit cell exhibits fourfold degeneracy at the symmetry points of K/K’ in its irreducible Brillouin zone (IRBZ). By combining the complementary parts, strong effective magnetoelectric coupling results in a nontrivial band gap at the K/K’ points. The topological structure is formed by aligning the unit cell with its flipped version, creating edge modes at the interface. The unidirectional propagation of line waves at the interface of two structural components is investigated in zig-zag and armchair configurations.
Using PTIs for leaky-wave antenna applications is not limited to spin PTIs[30, 31]. However, such structures require YIG photonic crystals and a static magnetic field bias, which limits their practicality. In contrast, there is no need for such requirements in spin PTI metasurfaces.
A leaky-wave antenna based on spin PTI and utilizing a zigzag arrangement has been investigated in [32]. However, this approach presents several significant challenges, including a narrow bandwidth, limited backward radiation scanning, coupling between radiating elements, and an inadequate radiation pattern.
@@
This paper proposes that adopting an armchair arrangement, instead of a zigzag design, effectively addresses the identified issues and facilitates backward-to-forward beam scanning. A crucial point is that broadside radiation in unbalanced composite right/left-handed (CRLH) leaky-wave antennas presents significant challenges, often leading to a sudden drop in gain due to an open stopband. However, in the proposed antenna with the armchair arrangement, the impact of this phenomenon is significantly reduced. Simulations were conducted using Ansys HFSS and were subsequently repeated in CST Studio to validate the results.
II Motivation to use armchair arrangement
The concept of utilizing the armchair arrangement as a leaky-wave antenna derives directly from the trapezoidal serpent line microstrip used in antenna applications[34]. As illustrated in Fig.1, there is a clear analogy between these two structures. Initially, to compare the zigzag and armchair configurations, the dimensions were selected to match those in [32]. Specifically, a periodic length of 7 mm and a border width of 0.25 mm were used. The substrate employed is Rogers/Duroid 5880, which has a thickness of 0.127 mm, a permittivity of 2.2, and a tangent loss factor of 0.0009. It is important to note that these dimensions are not final and may vary depending on the specific application. The parametric study on period length and border width identifies the dimensions necessary for designing an X-band leaky-wave.
@@ II Motivation to use armchair arrangement
II-A Parametric Study
The objective was to design an X-band leaky-wave antenna using a spin topological insulator featuring a hexagonal lattice arranged in an armchair configuration. Therefore, it is essential to analyze the cell parameter to determine the appropriate cell periodicity and border width. The Fig.3 illustrates how the limits of the bandgap vary with changes in the period length and border width. Specifically, Fig. 3 demonstrates that both the maximum and minimum bandgap will decrease as the period length increases. Furthermore, the decline rate of the upper limit (the minimum of the second degenerate band) is greater than that of the lower limit (the maximum of the first degenerate band). On the other hand, Fig.3 shows that the increasing of the border width will lead to a rise in the bandgap width, with a more significant rate of increase observed for shorter period lengths.
@@ II-A Parametric Study
II-B Matching ASL with Topological structure
@@
The field profile of the ASL and the identified cut planes are presented in Fig. 6, where the red and blue areas demonstrate the highest and the lowest amplitude of electric field,respectively. The direction of the triangles shows the direction of tangential electric field to the cut plane. All of these profiles exhibit confinement of the electric field in the interface of two side of structure but some of them have more similarity to the electric field of ASL cross section (see 6). Among these field profiles, those displayed in Figs. 6, 6 show lesser similarity to the ASL profile in Fig. 6 than the others, Because those cut lines are not perpendicular to the propagation direction.
@@
III Antenna application
The main advantage of using the armchair arrangement in antenna applications is that it effectively places most of the topological bandgap within the fast wave region. According to the findings in Appendix A, this design enables the proposed structure to scan through the broadside, a capability that conventional unbalanced CRLH leaky-wave antennas do not offer.
@@ III Antenna application
The variation of maximum realized gain with increasing frequency, as analyzed using Ansys HFSS and CST Studio, is illustrated in Fig. 13. The results are consistent and provide further validation of the simulations. The slight decline observed in the frequency range of 9.5 to 10 GHz is attributed to the inherent bandgap of the armchair arrangement. In contrast, to the best of our knowledge, other unbalanced CRLH leaky-wave antennas do not exhibit this characteristic [37, 2, 38, 39]. Another noteworthy feature is the nearly constant realized gain across most of the operational frequency range, which typically ranges from 4 to 6 dB.
IV Conclusions
A leaky-wave antenna is proposed using spin photonic topological insulators (PTIs). It features a hexagonal complementary unit cell arranged in an armchair configuration. The primary advantage of using the armchair arrangement in leaky-wave antenna applications is that most of topological bandgap is positioned within the fast wave region. In contrast, for the zigzag arrangement, most of the topological bandgap falls within the slow wave region. The design of the proposed antenna allows for a scanning range of 53 degrees, from 126 to 73 degrees, within an approximate bandwidth of 2.7 GHz. Importantly, the antenna sustains its performance at broadside angles without experiencing a significant decrease, even in the presence of an open stopband.
In Table I, a comparison of the proposed leaky-wave antennas utilizing topological insulators is presented. The table indicates that the antenna designed with an armchair arrangement offers the widest bandwidth and the highest scanning range. The HPMW of the proposed antenna is a weaker point compared to other references, particularly [33] and [30]. This limitation can be improved by increasing the antenna’s length. None of the references listed in the table can radiate in a broadside direction, except for [30]. However, this option has significant drawbacks, including a single beam pattern, a narrow bandwidth, a bulky design, and the requirement for a static magnetic field bias. In contrast, the proposed antenna not only radiates broadside but also offers two beams, a wide bandwidth, a low-profile design, and does not require any additional facilities.
@@ IV Conclusions
Appendix A The Zig-Zag and Armchair Arrangements Integration
@@
The realized gain pattern at 15.8 GHz, where radiation occurs in a broadside direction, is shown in Fig. 15. A comparison between the patterns of the zigzag arrangement (Fig. LABEL:zigzag) and the armchair arrangement reveals that the armchair structure is more suitable.
@@
At 15.8 GHz, the realized gain in the open stopband is less than -3 dB compared to the values observed at the ends of the scanning range. To the best of our knowledge, this unusual phenomenon does not occur in other unbalanced CRLH leaky-wave antennas. The 36-degree scan at both the top and bottom of the antenna surpasses the findings reported in [33], matching those in [32], even as it crosses the broadside.
Acknowledgment
We would like to thank Prof. A. Zeidaabadi Nezhad form Isfahan University of Technology and Dr. R.J.B. Davis from University of California, San Diego for the fruitful and illuminating discussions.